# Harvesting Private Medical Images in Federated Learning Systems with Crafted Models

### v1

@@
Abstract
Federated learning (FL) allows a set of clients to collaboratively train a machine-learning model without exposing local training samples. In this context, it is considered to be privacy-preserving and hence has been adopted by medical centers to train machine-learning models over private data. However, in this paper, we propose a novel attack named MediLeak that enables a malicious parameter server to recover high-fidelity patient images from the model updates uploaded by the clients. MediLeak requires the server to generate an adversarial model by adding a crafted module in front of the original model architecture. It is published to the clients in the regular FL training process and each client conducts local training on it to generate corresponding model updates. Then, based on the FL protocol, the model updates are sent back to the server and our proposed analytical method recovers private data from the parameter updates of the crafted module. We provide a comprehensive analysis for MediLeak and show that it can successfully break the state-of-the-art cryptographic secure aggregation protocols, designed to protect the FL systems from privacy inference attacks. We implement MediLeak on the MedMNIST and COVIDx CXR-4 datasets. The results show that MediLeak can nearly perfectly recover private images with high recovery rates and quantitative scores. We further perform downstream tasks such as disease classification with the recovered data, where our results show no significant performance degradation compared to using the original training samples.
Index Terms:
Federated Learning, Privacy Leakage, Medical Images.
I Introduction
@@
Federated learning (FL) has been regarded as a key enabling technology for the future implementation of AI-empowered medical diagnosis and treatment [1, 2, 3, 4, 5, 6, 7, 8]. FL allows medical centers to collaboratively train machine-learning models for various clinical tasks such as abnormality detection and disease classification, without sharing private patient information. This is very important because medical centers are bound to preserve patient privacy and their data usage is strictly restricted in many clinical applications. Under the FL framework, distributed training is set up in a way where hospitals that own private clinical data usually serve as clients, and a server – either hosted at one of the collaborating sites or maintained by a third party, integrates the model updates received from each client to orchestrate the federated learning paradigm. There are multiple open sources (such as NVFLARE [9] and OpenFL [10]) as well as commercial platforms (such as Rhino FCP [11]), designed to streamline the implementation of FL. During the FL training process, only the model updates, which refer to either the gradients or parameter updates, are exchanged between the participant and the server, and all the private training samples are kept locally at the clinical site. Therefore, when first introduced, federated learning was considered to be privacy-preserving and the model updates are regarded as safe vectors that hide training samples’ private information [12, 13, 14, 15].
@@
However, recent privacy attacks challenge this intuition as they demonstrate that a curious or malicious parameter server can extract useful information about the training samples such as their labels, membership information, and even training sample themselves from the model updates [16, 17, 18, 19, 20, 21, 22, 23, 24]. Of particular interest, the model inversion attacks (MIAs) [20, 22, 23, 24, 25] are a type of privacy attack aiming to recover the original training images. They take the individual model updates provided by the clients as inputs and reverse them back to the local training images. This could be detrimental in radiologic applications, where such an attack can reconstruct patient-specific data. Existing MIAs usually formulate this reverse process as an optimization problem and can achieve good recovery performance to recover high-fidelity training images from the model updates, when undergoing enough optimization iterations, which completely exposes the information that the FL system wants to protect. Fortunately, as the arms race between the attacker and defender goes on, a specialized multi-party computation (MPC) mechanism named secure aggregation (SA) protocol is proposed to prevent MIAs [26, 27, 28]. The fundamental idea of SA protocols is to use various cryptography primitives (e.g., secret sharing) to mask the individual model updates with random values but keep their summation identical to the pre-masked value so that the FL system can proceed to the training process without exposing individual model updates. As a result, the SA protocol prevents the MIAs from even obtaining the inputs (individual model updates) to the reverse processes, and the privacy of the FL systems is guaranteed. Moreover, existing MIAs are also facing scalability and efficiency challenges, as in practice they need the server to consume extensive resources (usually hundreds of seconds) to recover only a few images, making them hardly employable in real systems.
@@
In this paper, we propose a novel attack name MediLeak that addresses the limitation of the existing works, making it a very practical attack and representing a privacy vulnerability of the FL systems. MediLeak can recover hundreds of training samples in a batch from the victim client simultaneously within a few seconds. MediLeak can also break the secure aggregation protocols as it can recover the training samples directly from the summation value of the model updates even though the individual ones are cryptographically masked. Technically, MediLeak is a two-phase attack including the attack preparation phase and the sample recovery phase. In the first phase, the attacker adds an additional two-linear layer module in front of the original model architecture and initializes the module with customized parameters before sending it to the clients. For the target victim, the attacker initializes the parameters of the two-layer module to form a “linear leakage” module with the help of an auxiliary dataset that has the same data format and distribution as the training samples. This “linear leakage” module is a powerful mathematical tool that can perfectly reverse its gradients back to its inputs, which are identical to the training samples because we place this module as the first component of the model architecture. For other clients, their two-layer modules’ parameters are crafted to form a “zero gradient” module, aiming to zero out their gradients and model updates. By doing so the aggregated model update is identical to the model update of the victim because all others are set to zero, rendering the SA protocols useless.
@@
We implement MediLeak on the MedMNIST [29] and COVIDx CXR-4 [30] datasets and evaluate our attack performance using the recovery rate, structural similarity (SSIM) score, peak signal-to-noise ratio (PSNR) score, and attack time. The results show that MediLeak achieves excellent performance on both datasets to recover hundreds of images simultaneously with high recovery rates and quantitative scores, with only a few seconds of execution time. We visualize the recovered images and they are visually indistinguishable from the original images. We compare the performance of MediLeak with three existing MIAs. The results demonstrate that MediLeak achieves better quantitative scores and is much more efficient than them. We further feed the recovered images to downstream disease diagnosis tasks. Our results show that the recovered images achieve classification performance close to the original ones, demonstrating the effectiveness of our attack.
In summary, this paper makes the following contributions:
  1. 1.
@@
    We propose MediLeak, a novel and powerful MIA that can recover large-batch and high-fidelity local training samples owned by the clients from the model updates efficiently when assuming the parameter server is a proactive attacker curious about the client’s private information, even when the system is protected by the state-of-the-art cryptography-based defense mechanisms.
  2. 2.
    Our attack represents a fundamental and practical privacy vulnerability of the current federated learning system as it leaks its participants’ private information, and promotes the urgent need for more effective defense mechanisms against such a powerful privacy attack.
  3. 3.
@@
    We provide rigorous mathematical analysis and proof for our attack. Our attack is a closed-form attack and does not incur any computation-intensive optimization process, which reduces the computational costs significantly compared to existing model inversion attacks.
  4. 4.
@@
    We implement MediLeak on medical image datasets under different practical assumptions in the FL systems. The results show that our attack can nearly perfectly recover the local training samples of the target victim with high performance in downstream clinical tasks.
II Background
II-A Federated Learning
@@
We consider for each training round $t$, there are $n$ clients denoted by $\mathcal{C}=\{c_{1},c_{2},\cdots,c_{n}\}$ to be selected by the parameter server $S$ to collaboratively train a global model $G=f_{\theta}:\mathcal{X\to Y}$, with each client $c_{i}$ holds a local dataset $D_{i}$. In detail, the parameter server $S$ first publishes the global model parameter $\theta^{t}$ to the clients. Then each client trains the received global model $G_{t}$ for $L^{t}_{i}$ local rounds over $D_{i}$ to produce its model update $\delta_{i}^{t}$. Note that when $L^{t}_{i}=1$, the model update $\delta_{i}^{t}$ can be replaced by the gradient $g_{i}^{t}$. After that, the client $c_{i}$ sends the model update $\delta_{i}^{t}$ back to the server $S$ and the server employs the federated average (FedAVG) algorithm [12] to conduct the training process:
$\theta^{t+1}=\sum_{i=1}^{n}\alpha_{i}\delta^{t}_{i},$ (1)
@@
where $\alpha_{i}$ is the weight assigned to client $c_{i}$. The summation of all weights $\{\alpha_{i}\}_{i:c_{i}\in\mathcal{C}}$ is 1 and can be adjusted according to the size of local datasets $D_{i}^{t}$ to avoid training bias. The server may also employ the FedSGD algorithm [12] as an alternative to conducting the training process. In the following sections, we will omit the notation $t$, because our attack and analysis are all conducted in a single FL training round.
II-B Model Inversion Attacks
@@
The model inversion attacks (MIAs) take the individual model updates $\delta_{i}$ as the inputs and aim to reconstruct them back to the local datasets $D_{i}$ held by the clients. This reversion problem has been formalized as an optimization problem, represented as $\arg\min_{\hat{D_{i}}}[d(\nabla\hat{D_{i}}-\nabla D_{i})]$, where $\hat{D_{i}}$ refer to randomly initialized dummy samples and $d()$ refers to a distance function such as the second norm distance. To solve this optimization problem, [20] utilizes the L-BFGS optimizer to reconstruct the dummy samples iteratively step by step until reaching a good optimization point. It is further improved by an analytical method that aims to recover the ground-truth labels of dummy samples from the gradients [21], which significantly eases the optimization task and helps accomplish better attack performance. Later works improve the optimization tools and focus on recovering larger batches of images on more practical machine learning models such as the ResNet [22, 23, 31, 32]. However, their recovery sizes are still restricted to the scale of tens, representing a scalability challenge. Moreover, all the aforementioned MIAs require costly iterative optimization methods in their design, which unfortunately, introduce a very large overhead to recover each batch of input images. These methods also cannot bypass the current SA protocols.
Figure 1: Threat Model. The server is considered to be a malicious attacker. The secure aggregation protocol is considered to be in place to protect the individual model updates.
Figure 2: MediLeak attack flow. MediLeak is a two-phase attack. In the first preparation phase, the attacker generates the adversarial global model. In the second reconstruction phase, the attacker sends the adversarial models to the clients and recovers the local samples when it receives their feedback.
II-C Secure Aggregation
@@
To enhance the privacy of the federated learning systems, Bonawitz et al. [26] propose a new type of specialized MPC mechanism named the secure aggregation (SA) protocols to fulfill an abstract function of masking individual model updates $\delta_{i}$ with random bits while keeping the summation of the masked values $\sum_{i=1}^{n}u_{i}$ identical to the pre-masked one $\sum_{i=1}^{n}\delta_{i}$. Therefore, despite variations of detailed cryptographic design, all SA protocols ensure that the server cannot obtain the individual model updates $\delta_{i}$ to launch any model inversion attacks, but can proceed with the FL training process with the aggregated model update $\sum_{i=1}^{n}u_{i}$, which is identical to $\sum_{i=1}^{n}\delta_{i}$. Since its initial introduction, the SA protocols have been continuously refined to incorporate other properties including communication efficiency, drop-out resilience [33, 34, 35, 36], and security against malicious clients [27, 28], making it the current state-of-the-art privacy protection mechanism for FL systems.
@@
Secure Aggregation under Challenge: Under the honest-but-curious attack model, the SA protocols have proven to be secure against various MIAs. However, recent works adopt a stronger attack model to assume a proactive attacker that modifies the global model’s parameters and even its architecture before publishing it to the clients. Under this assumption, [25] proposes a novel attack that retrieves a target individual model update from the aggregated result that breaks the SA protocols. The fundamental idea of this novel attack is to craft the model parameters to adversarial models and distribute different adversarial models to different clients strategically. The adversarial models are crafted to ensure that only the model update of the victim client is preserved while all the others are zeroed out. The limitation of this attack is that it incorporates a costly optimization process in its design, which introduces too much attack overheads. [24] proposes another attack method to add crafted modules before the original model architecture. These additional modules are crafted with delicate mathematical designs to ensure that the model gradients can be perfectly reversed back to inputs whenever the server receives any model updates. The limitation of this attack is that the attacker cannot link the recovered images to their owners, which means the SA protocols still preserve a certain level of privacy, known as “privacy by shuffling”.
In summary, the security of the SA protocols is challenged by the advanced attacks under the proactive attacker assumption. In our design, we adopt this fundamental attack assumption to break the SA protocol and address various limitations of the existing work at the same time.
III Threat Model

### v2

# MedLeak: Multimodal Medical Data Leakage in Secure Federated Learning with Crafted Models

@@
Abstract.
Federated learning (FL) allows participants to collaboratively train machine learning models while keeping their data private, making it ideal for collaborations among healthcare institutions on sensitive datasets. However, in this paper, we demonstrate a novel privacy attack called MedLeak, which allows a malicious participant who initiates the FL task as the server to recover high-quality site-specific private medical images and text records from the model updates uploaded by clients. In MedLeak, a malicious server introduces an adversarially crafted model during the FL training process. Honest clients, unaware of the insidious changes in the published model, continue to send back their updates as per the standard FL training protocol. Leveraging a novel analytical method, MedLeak can efficiently recover private client data from the aggregated parameter updates. This recovery scheme is significantly more efficient than the state-of-the-art solutions, as it avoids the costly optimization process. Additionally, the scheme relies solely on the aggregated updates, thus rendering secure aggregation protocols ineffective, as they depend on the randomization of intermediate results for security while leaving the final aggregated results unaltered.
We implement MedLeak on medical image datasets MedMNIST, COVIDx CXR-4, and Kaggle Brain Tumor MRI datasets, as well as the medical text dataset MedAbstract. The results demonstrate that the proposed privacy attack is highly effective on both image and text datasets, achieving high recovery rates and strong quantitative scores. We also thoroughly evaluate MedLeak across different attack parameters, providing insights into key factors that influence attack performance and potential defenses. Furthermore, we perform downstream tasks, such as disease classification, using the recovered data, showing no significant performance degradation compared to the original training samples. Our findings validate the need for enhanced privacy measures in federated learning systems, particularly for safeguarding sensitive medical data against powerful model inversion attacks.
Federated Learning, Model Inversion Attack, Medical AI Privacy.
journalyear: 2025copyright: ccconference: ACM/IEEE International Conference on Connected Health: Applications, Systems and Engineering Technologies; June 24–26, 2025; New York, NY, USAbooktitle: ACM/IEEE International Conference on Connected Health: Applications, Systems and Engineering Technologies (CHASE ’25), June 24–26, 2025, New York, NY, USAdoi: 10.1145/3721201.3721375isbn: 979-8-4007-1539-6/2025/06ccs: Security and privacyccs: Computing methodologies Machine learningccs: Security and privacy
1. Introduction
@@
Federated learning (FL) has developed as a key enabling technology for the future implementation of AI-powered medical diagnosis and treatment systems (Brisimi et al., 2018; Rieke et al., 2020; Sheller et al., 2020; Adnan et al., 2022; Pati et al., 2021; Dayan et al., 2021; Nguyen et al., 2022; Ng et al., 2021). FL allows medical centers to collaboratively train machine-learning models for various clinical tasks such as disease classification and clinical diagnosis, without sharing private patient information. This is crucial because medical centers are bound to preserve patient privacy and their data usage is strictly restricted in many clinical applications in accordance with regulatory guidelines. Under the FL framework, distributed training is set up in a way where hospitals that own private clinical data usually serve as clients, and a server – either hosted at one of the collaborating sites or maintained by a third party, integrates the model updates received from each client to orchestrate the federated learning paradigm. There are multiple open source (such as NVFLARE (Roth et al., 2022) and OpenFL (Reina et al., 2021)) as well as commercial platforms (such as Rhino FCP (Stephens, 2021)), designed to streamline the implementation of FL. During the FL training process, only the model updates, which refer to either the gradients or parameter updates, are exchanged between the participant and the server, while the private training samples are kept securely at the clinical site. Therefore, when first introduced, federated learning was considered to be privacy-preserving and the model updates are regarded as safe vectors that hide training samples’ private information (McMahan et al., 2017; Zhang et al., 2021; Wei et al., 2021; Pfitzner et al., 2021; Wang et al., 2022).
@@
Recent privacy attacks challenged the privacy-preserving property associated with FL. It is demonstrated that a curious or malicious parameter server can extract information related to the training samples such as their labels, membership information, and even the whole training sample using the model updates (Nasr et al., 2019; Fu et al., 2022; Luo et al., 2021; Wang et al., 2019; Zhu et al., 2019; Zhao et al., 2020; Geiping et al., 2020; Yin et al., 2021; Fowl et al., 2021). Of particular interest, the model inversion attacks (MIAs) (Zhu et al., 2019; Geiping et al., 2020; Yin et al., 2021; Fowl et al., 2021; Pasquini et al., 2022; Shi et al., 2025) are a type of privacy attack aiming to recover the original training images. They take the individual model updates provided by the clients as inputs and reverse them back to the local training samples. This could be detrimental in medical applications, where such an attack can reconstruct patient-specific data. Existing MIAs usually formulate this reverse process as an optimization problem and have been shown to achieve good recovery performance in recovering high-fidelity training images from the model updates, when enough optimization iterations are performed. This completely exposes the information that the FL system has been designed to protect.
However, existing optimization-based MIAs are facing serious scalability and efficiency challenges. In practice, they need the server to consume extensive resources (usually hundreds of computation seconds with large memory) to recover only a few images, making them virtually impractical for real-world systems. Further, such MIAs can also be prevented by a specialized multi-party computation (MPC) mechanism named secure aggregation (SA) protocol (Bonawitz et al., 2017; Pillutla et al., 2022; Burkhalter et al., 2021), whose fundamental idea is to use various cryptography primitives (e.g., secret sharing) to mask individual model updates with random values but keep their summation identical to the pre-masked value. In this way, the FL system can proceed to the training process without exposing individual model updates, preventing the MIAs from reversing them back to local training samples.
@@
In this paper, we propose a novel attack name MedLeak that addresses the limitation of the existing works. Our attack design makes it very practical, which could be detrimental to privacy in existing FL systems. MedLeak is an efficient attack, capable of recovering hundreds of training samples in a batch from the victim client within just a few seconds. MedLeak can also break the SA protocols as it can recover the training samples directly from the summation value of the model updates even though the individual ones are cryptographically masked. In addition, MedLeak can be accomplished within one FL training round, making it very practical and stealthy. To further evade detection, the attacker can launch the attack in the initial FL training rounds to avoid hurting the FL training performance. A greedy attacker can even launch the attack multiple times during continuous training rounds to harvest as many local training samples as possible.
@@
Technically, MedLeak is a two-phase attack including the attack preparation phase and the sample recovery phase. In the first phase, the attacker adds an additional two-linear layer module in front of the original model architecture and initializes the module with customized parameters before sending it to the clients. For the target victim, the attacker initializes the parameters of the two-layer module to form a “linear leakage” module with the help of an auxiliary dataset that has the same data format and distribution as the training samples. This “linear leakage” module is a powerful mathematical tool that can perfectly reverse its gradients back to its inputs, which are identical to the training samples because we place this module as the first component of the model architecture. For other clients, their two-layer modules’ parameters are crafted to form a “zero gradient” module, aiming to zero out their gradients and model updates. By doing so the aggregated model update is identical to the model update of the victim because all the others are set to zero, rendering the SA protocols useless.
As medical data is usually composed of both image and text records, we explore and extend MedLeak’s capability to recover both data modalities. Particularly, the recovery of medical text records is largely ignored in the existing literature and we are proposing the first work in this direction to our knowledge. Compared to medical images, the recovery of medical text records is more challenging because they are discrete natural language words with different paragraph lengths in the input space, resulting in completely different FL model architectures and parameter calculation methodologies. To address this challenge, we slightly modify and customize our attack to insert the malicious module after the embedding layer rather than at the front of the whole model to first launch the recovery word embedding vectors, and then further reverse these word embedding vectors back to input tokens (words).
Table 1. A comparison between the existing MIAs and MedLeak with respect to their attack assumptions, efficiency, scales, capabilities, and attack generalizability.
Attacks Attack Model Attack Efficiency Scale Break SA? Recover Text? Targeted?
(Zhu et al., 2019; Zhao et al., 2020; Geiping et al., 2020; Yin et al., 2021; Lu et al., 2022; Wen et al., 2022) Honest-but-curious Low (Optimization-based) $10^{1}$ No No (Image-only) No
(Pasquini et al., 2022) Malicious server Low (Optimization-based) $10^{1}$ Yes No (Image-only) Yes
(Fowl et al., 2021; Zhao et al., 2023) Extra attack module High (Mostly closed-form) $10^{2}$ to $10^{3}$ Yes No (Image-only) No
MedLeak Extra attack module High (Closed-form) $10^{2}$ to $10^{3}$ Yes Yes Yes
@@
We evaluated MedLeak on MedMNIST (Yang et al., 2023), COVIDx CXR-4 (Wang et al., 2020), Kaggle Brain Tumor MRI (Nickparvar, 2022) datasets for medical images, and on the MedAbstract dataset (Schopf et al., 2023) for medical text records. For medical images, we evaluated our attack performance using the recovery rate, structural similarity (SSIM) score, peak signal-to-noise ratio (PSNR) score, and attack time. Our results show that MedLeak achieves excellent performance on these datasets to recover hundreds of images simultaneously with high recovery rates and quantitative scores, with only a few seconds of execution time. On visual inspection, the recovered images are virtually indistinguishable from the original images. We compared the performance of MedLeak with three existing MIAs. Our results demonstrate that MedLeak achieves better quantitative scores and is comparatively much more efficient. We further fed the recovered images to downstream disease diagnosis tasks. Our results show that the recovered images achieved a classification performance close to the original images, validating the effectiveness of our attack. For medical text records, we evaluated our attack performance using the recovery rate, word error rate (WER), and attack time. The results show that MedLeak can accurately recover tens of long clinical data paragraphs (e.g., descriptions of patients’ health conditions) spanning up to hundreds of words simultaneously. This highlights MedLeak’s strong capability to recover non-image modality records effectively and efficiently.
In summary, in this paper we present the following contributions:
  1. (1)
@@
    We propose MedLeak, a novel and powerful MIA that is capable of recovering high-quality local training samples in large batches using model updates from FL clients efficiently, even when state-of-the-art cryptography-based defense mechanisms such as secure aggregation are employed.
  2. (2)
    Our attack represents a fundamental and practical privacy vulnerability of the medical FL system as it compromises individual clients’ privacy. MedLeak can target both medical image and text data, demonstrating its broad applicability in the medical domain.
  3. (3)
@@
    We provide rigorous mathematical analysis and proof for our attack. Our proposed attack design is closed-form, hence avoids incurring any computation-intensive optimization and significantly reduces the computational costs when compared to existing MIAs.
  4. (4)
@@
    We implement MedLeak on medical images and text datasets under different practical assumptions in the FL systems. The results show that our attack can nearly perfectly recover different types of local training samples of the target victim.
2. Background
2.1. Federated Learning
@@
We consider for each training round $t$, there are $n$ clients denoted by $\mathcal{C}=\{c_{1},c_{2},\cdots,c_{n}\}$ to be selected by the parameter server $S$ to collaboratively train a global model $G=f_{\theta}:\mathcal{X\to Y}$, with each client $c_{i}$ holds a local dataset $D_{i}$. In detail, the parameter server $S$ first publishes the global model parameters $\theta^{t}$ to the clients. Then each client trains the received global model $G_{t}$ for $L^{t}_{i}$ local rounds over $D_{i}$ to generate its model update $\delta_{i}^{t}$. Note that when $L^{t}_{i}=1$, the model update $\delta_{i}^{t}$ can be replaced by the gradient $g_{i}^{t}$. After that, the client $c_{i}$ sends the model update $\delta_{i}^{t}$ back to the server $S$ and the server employs model aggregation (such as using the federated average (FedAVG) algorithm (McMahan et al., 2017)) to conduct the training process:
(1) $\theta^{t+1}=\sum_{i=1}^{n}\alpha_{i}\delta^{t}_{i},$
@@
where $\alpha_{i}$ is the weight assigned to client $c_{i}$. The summation of all weights $\{\alpha_{i}\}_{i:c_{i}\in\mathcal{C}}$ is 1 and can be adjusted according to the size of local datasets $D_{i}^{t}$ to avoid training bias. The server may also employ alternative aggregation strategies (such as FedSGD algorithm (McMahan et al., 2017)) for model aggregation. In the following text, we will omit the notation $t$, because our attack and analysis are all conducted in a single FL training round.
Federated learning has been widely used in the healthcare domain, enabling different healthcare providers to collaborate with each other towards accomplishing the training of machine learning models for both natural language processing and imaging tasks (Pfitzner et al., 2021; Lin et al., 2021; Liu et al., 2021). Such clinical FL systems usually leverage the patients’ private information as the local training datasets. These can be the patients’ radiology scans, textual reports, and tabular records describing the patients’ visits and health conditions. This information is considered to be highly private and sensitive and is protected by strict governance laws such as HIPPA and GDPR. The medical FL system is deployed to protect such privacy by ensuring that the sensitive data never leaves the firewalls within the healthcare sites during the model training process. However, herein we will demonstrate with MedLeak that the privacy-preserving property of current FL systems is under challenge.
2.2. Model Inversion Attacks
@@
The model inversion attacks (MIAs) take the individual model updates $\delta_{i}$ as the inputs and aim to reconstruct them back to the local datasets $D_{i}$ held by the clients. This reversion problem has been formalized as an optimization problem, represented as $\arg\min_{\hat{D_{i}}}[d(\nabla\hat{D_{i}}-\nabla D_{i})]$, where $\hat{D_{i}}$ refer to randomly initialized dummy samples and $d()$ refers to a distance function such as the second norm distance. To solve this optimization problem, (Zhu et al., 2019) utilizes the L-BFGS optimizer to reconstruct the dummy samples iteratively step by step until reaching a good optimization point. It is further improved by an analytical method that aims to recover the ground-truth labels of dummy samples from the gradients (Zhao et al., 2020), which significantly eases the optimization task and helps accomplish better attack performance. Later works have improved the optimization tools and focus on recovering larger batches of images on more practical machine learning models such as the ResNet (Geiping et al., 2020; Yin et al., 2021; Lu et al., 2022; Wen et al., 2022). However, their recovery sizes are still restricted to the scale of tens, representing a scalability challenge. Moreover, all the aforementioned MIAs require costly iterative optimization methods in their design, which introduce a very large overhead to recover each batch of input images. Further, these existing methods also cannot bypass the current SA protocols.
2.3. Secure Aggregation
@@
To enhance the privacy of the federated learning systems, Bonawitz et al. (Bonawitz et al., 2017) proposed a new type of specialized MPC mechanism named the secure aggregation (SA) protocols to fulfill an abstract function of masking individual model updates $\delta_{i}$ to $u_{i}$ with random bits, while keeping the summation of the masked values $\sum_{i=1}^{n}u_{i}$ identical to those of the pre-masked values $\sum_{i=1}^{n}\delta_{i}$. Therefore, despite variations of detailed cryptographic design, all SA protocols ensure that the server cannot obtain the individual model updates $\delta_{i}$ to launch any model inversion attacks, but can proceed with the FL training process with the aggregated model update $\sum_{i=1}^{n}u_{i}$, which is identical to $\sum_{i=1}^{n}\delta_{i}$. Since its initial introduction, the SA protocols have been continuously refined to incorporate other properties including communication efficiency, drop-out resilience (Bell et al., 2020; Choi et al., 2020; Guo et al., 2020; Kadhe et al., 2020), and security against malicious clients (Pillutla et al., 2022; Burkhalter et al., 2021), making it the current state-of-the-art privacy protection mechanism for FL systems.
@@
Secure Aggregation under Challenge: Under the honest-but-curious attack model, the SA protocols have proven to be secure against various MIAs. However, recent works adopt a stronger attack model to assume a proactive attacker that modifies the global model’s parameters and even its architecture before publishing it to the clients. Under this assumption, (Pasquini et al., 2022) proposed a novel attack that retrieves a target individual model update from the aggregated result, breaking the SA protocols. The fundamental idea is to craft the model parameters to adversarial models and distribute different adversarial models to different clients strategically. The adversarial models are crafted to ensure that only the model update of the victim client is preserved while all the others are zeroed out. The limitation of this attack is that it incorporates a costly optimization process in its design, which introduces too much attack overheads. (Fowl et al., 2021) proposed another attack method to add crafted modules before the original model architecture. These additional modules are crafted with delicate mathematical designs to ensure that the model gradients can be perfectly reversed back to inputs whenever the server receives any model updates. The limitation of this attack is that the attacker cannot link the recovered images to their owners, which means the SA protocols still preserve a certain level of privacy, known as “privacy by shuffling”. Later, (Zhao et al., 2023) addressed this issue by designing a more complex adversarial module composed of convolutional and linear layers before the original model to identify the client associated with the recovered images. However, using a computer vision-related architecture prohibits the attack from being adopted to recover text records.
2.4. Attack Summary and Comparison
In Tab. 1, we summarize and compare the existing MIAs according to their attack assumptions, efficiency, scales, capabilities, and generability. In our design, we adopt the active attacker assumption to break the SA protocol and simultaneously address various limitations of the existing works. In particular, we aim to address the following three problems: 1) attack efficiency-our attack shall not employ any computation-intensive optimization process to incur large overheads; 2) attack effectiveness- our attack’s capability of recovering high-quality local training samples from the aggregated model updates and attribute them back to individual clients even when SA protocols are incorporated, and 3) attack generality- our attack can be applied to traditional image recovery tasks as well as text recovery task (currently less explored).
3. Threat Model
Figure 1. Threat Model. The server is considered to be a malicious attacker. The secure aggregation protocol is considered to be in place to protect the individual model updates.
Figure 2. MedLeak attack flow. MedLeak is a two-phase attack. In the first preparation phase, the attacker generates the adversarial global model. In the second reconstruction phase, the attacker sends the adversarial models to the clients and recovers the local samples when it receives their feedback. MedLeak can reconstruct both image and non-image data and this figure demonstrates the reconstruction of the medical radiology images.